A consultant charges $45 for each hour she works on a consultation, plus a flat $30 consulting fee. How many hours of work are included in a $210 bill for a
consultation?
a)2 4/5 b) 4 c) 4 2/3 d) 5 1/2 e) 7

Answer by richwmiller(17219)

45x+30=210
45x=180
x=4 hours
check
4*45+30=210
180+30=210
210=210
ok